Choose the system of equations in x and y describing the situation.
Leroy has $15 less in his pocket than Donna. If Donna hs $8 more, she would have 4 times as much money as Leroy. This confuses me. Which do I use to figure out the problem?
a) x=y-15
y-8=4x
b) x=y-15
y+8=4x
c) x=y-15
y+8=x+4
d) y=x-15
y+8=4x

Answer by alicealc(293)

Leroy's money = x
Donna's money = y

Leroy has $15 less in his pocket than Donna:
x = y - 15

If Donna has $8 more, she would have 4 times as much money as Leroy:
y + 8 = 4x

so, the answer is
b) x = y - 15
y + 8 = 4x